OpenStudy (anonymous):

Solve t = the quantity p plus 3 r, divided by six for r

OpenStudy (anonymous):

t = (p+3r)/6 starting equation 6t = p+3r multiply both sides by six 6t - p = 3r subtract p from both sides (6t - p)/3 = r divide both sides by 3